Circular No. 8896

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V597 PUPPIS = NOVA PUPPIS 2007 N. N. Samus, Institute of Astronomy, Russian Academy of Sciences,

informs us that this nova (cf. IAUC 8895) has been given the designation V597 Pup. A. J. S. Pereira, Carnaxide, Portugal, writes that the nova was at visual mag 7.4 on Nov. 15.19 and 15.24 UT.

H. Naito and N. Tokimasa, Nishi-Harima Astronomical Observatory, report that they obtained a low-resolution spectrum (range 410–670 nm; resolution 1200 at Hα) of V597 Pup on Nov. 14.77 UT with the 2.0-m NAYUTA telescope (+ MALLS). The spectrum shows broad Balmer lines and Fe II (42 multiplet) with P-Cyg features, as well as blueshifted absorption lines of Fe II (49, 74), O I 615.5–615.8 nm, and Si II 634.7-nm. The nova belongs to the Fe II type and was caught near maximum light. A narrow Na I D absorption line (EW = 0.3 nm) indicates that some interstellar reddening exists. The expansion velocities derived from the FWHM of the Balmer features Hα, Hβ, and Hγ are 1800, 1700, and 1800 km/s, respectively.

COMETS C/2007 R9, C/2007 R10, C/2007 S3, C/2007 S4 (SOHO) Further to IAUC 8895, additional near-sun comets have been found

on SOHO website images, with their “discovery” observations tabulated below; all were Kreutz sungrazers except for C/2007 R10 (Meyer group). C/2007 R9 and C/2007 S3 were tiny and stellar in appearance (mag ∼ 7). C/2007 R10 was also stellar in appearance, very small, and faint (mag ∼ 7.5). C/2007 S4 appeared quite diffuse and very faint (mag ∼ 7) in C3 images, and extremely faint and diffuse in C2 images.

Comet 2007 UT α_{2000} δ_{2000} Inst. F MPEC
C/2007 R9 Sept. 13.446 11^{h}07.^{m}8 + 3^{o}02′ C3  BZ 2007-U36
C/2007 R10  14.938 11 25.8 + 5 13 C2  AK 2007-U36
C/2007 S3  16.113 11 17.8 + 1 59 C3  BZ 2007-U36
C/2007 S4  23.763 11 47.7 − 1 38 C3/2 MU 2007-U36

COMET 17P/HOLMES Naked-eye total-magnitude and coma-diameter estimates: Nov. 6.83

UT, 2.6, 22′ (T. Karhula, Vasteras, Sweden); 11.69, 2.7, 33′ (S. Yoshida, Kanagawa, Japan); 12.81, 2.7, 35′ (M. Meyer, Limburg, Germany); 13.62, 2.9, 30′ (S. Yoshida, Kanagawa, Japan); 14.03, 2.6, 30′ (J. J. Gonzalez, Leon, Spain); 15.13, 2.9, ∼ 32′ (A. Pereira, Cabo da Roca, Portugal).

2007 November 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green